need steps on this algebra problem -7x+2=54-3
Answer by addingup(3677)

7x+2=54-3
7x+2=51
7x= 49 now divide both sides by 7, and you get:
x= 7
Proof:
7(7)+2= 54-3
49+2= 51
51= 51 Our answer is correct